# MS-HuBERT: Mitigating Pre-training and Inference Mismatch in Masked Language Modelling methods for learning Speech Representations

### v1

@@ Abstract
In recent years, self-supervised pre-training methods have gained significant traction in learning high-level information from raw speech. Among these methods, HuBERT has demonstrated SOTA performance in automatic speech recognition (ASR). However, HuBERT’s performance lags behind data2vec due to disparities in pre-training strategies. In this paper, we propose (i) a Swap method to address pre-training and inference mismatch observed in HuBERT and (ii) incorporates Multicluster masked prediction loss for more effective utilization of the models capacity. The resulting method is, MS-HuBERT, an end-to-end self-supervised pre-training method for learning robust speech representations. It beats vanilla HuBERT on the ASR Librispeech benchmark on average by a 5% margin when evaluated on different finetuning splits. Additionally, we demonstrate that the learned embeddings obtained during pre-training encode essential information for improving performance of content based tasks such as ASR.
@@ 1 Introduction
In the recent years, there has been a significant interest in studying self-supervised pre-training methods to learn/encode high level information present in the speech data [1, 2, 3, 4, 5]. These SSL methods utilize the input data itself to learn to encode useful information, with the choice of pretext task playing a pivotal role in the encoded information. The most popular pretext task used is masked predictive coding (MPC) [6, 1, 7, 8, 9]. HuBERT [2] is one such model that popularized the masked language modelling (MLM) technique to learn high-level speech representations from raw audio by achieving state-of-the-art (SOTA) on the automatic speech recognition (ASR) task. The underlying concept of HuBERT revolves around iterative pre-training: starting with a raw audio/pseudo-label pair (x/y), the model undergoes successive training iterations where the trained model updates the pseudo-labels, iteratively refining its representations until a predefined stopping criterion is reached. However, despite its success, HuBERT falls short compared to data2vec [10] in ASR performance for two primary reasons: firstly, during pre-training, data2vec accesses the full context to generate continuous labels, which are updated after each gradient update step, as opposed to the fixed discrete labels utilized in HuBERT for the each iteration; and secondly, the ground truth labels are created by averaging the representations from multiple layers for loss calculation.
@@
To bridge this gap, we propose two modifications to the HuBERT framework. Firstly, we introduce the ”Swap” method to enable full context access during pre-training, thus addressing the pre-training and inference mismatch observed in HuBERT and other MLM-based methods by using both the masked and unmasked views during pre-training. Swap is motivated by a simple idea, used heavily in the field of computer vision [11, 12, 13]. Where two augmented views of the input are used to learn a high level representation. Given $e$ layers in a encoder, it is a general practice to add a similarity loss on the output embeddings of the encoder after each layer, as seen in works using U-net type architectures [14, 15, 16]. In contrast, our proposed Swap method simply swaps the output embedding, at certain indices, after each layer of the encoder between the masked and unmasked view of the input. Motivated by the simple fact, that the learned model is expected to generate exactly the same output regardless of the two views, Swap method guides the output after each layer to achieve the same.
@@
Secondly, inspired by the work of Yadav et al. [17], we incorporate a Multicluster masked prediction loss (MPL) approach. Using multiple cluster centers, also called multiple resolutions, has been investigated by [18, 19, 17]. In [19], the author introduces down-sampling and up-sampling modules within the transformer encoder after each layer to facilitate learning features at multiple resolutions. On the other hand, [18] explores parallel and hierarchical variations of HuBERT with findings indicating the superiority of the hierarchical approach. This involves training multiple models, each model adds almost same parameters as the original HuBERT, at various resolutions using CNN as a down-sampling module. Lastly [17] uses the fact that MPL is applied at multiple layers of encoder at different resolutions. This method does not introduce any additional parameters to the original HuBERT model, except the linear layers used for loss calculation which are discarded after the pre-training. In this work, we adopt this approach and modify it for our use case for loss calculation given its simplicity and higher performance.
@@
Based on these observations. In this work, we propose MulticlusterSwap-HuBERT (MS-HuBERT) method, incorporates (i) our proposed Swap method to address the pre-training and inference mismatch issue, as the [MASK] symbol never appears during the inference, and (ii) the Multicluster MPL similar to [17]. Our contributions are as follows:
@@
    We show that MS-HuBERT outperforms the original HuBERT on the ASR Librispeech benchmark by a large margin. And matches the performance of data2vec in high-resource setting.
@@
    We showcase that the embeddings acquired during pre-training encode crucial information essential for addressing content based tasks such has ASR and phoneme recognition (PR). This shows the effective utliziation of the modeling capacity.
@@ 4 Results
4.1 Main Results: Supervised Fine-tuning and Inference
@@
Notably, upon the removal of the Swap concept, we observed a degradation in performance, particularly in low-resource settings. This proves that the Swap method does indeed contribute positively to the performance gains particularly in low resource setting.
@@ 4.2 MS-HuBERT as a Feature Extractor
Table 2: SUPERB fine-tuning results.
Method PR ASR
wav2vec 2.0 [1] 5.74 6.43
HuBERT [2] 5.41 6.42
WavLM [9] 4.84 6.21
data2vec [10] 4.69 4.94
MS-HuBERT 4.42 5.60
@@
Figure 2: Solid lines show the CCA similarity with the word labels. Dotted lines show the AUC area under the curve for them respectively.
CCA Similarity with Word Labels: Following the layer-wise analysis conducted by Pasad et al. [24, 25], we use a modified version of canonical correlation analysis (CCA). Specifically, a projection-weighted CCA (PWCCA) [26]. The plots are shown in Figure 2. It is clear that MS-HuBERT significantly enhances the performance of word-level information across the transformer encoder layers. Additionally, we compute the area under the curve (AUC) and observe that it consistently surpasses that of Hubert. This increases the model capacity utilization compared to HuBERT. We also observe that Swap increases the CCA similarity in the later layers, which could be the reason of performance gain in low resource settings on ASR.
@@ 5 Discussion
Given MS-HuBERT is trained using the pseudo labels generated from the the first iteration HuBERT, we posit that training two iterations using the MS-HuBERT methodology could potentially surpass data2vec on the ASR Librispeech benchmark and ASR tasks within the SUPERB benchmark.
Complexity and Computational Cost: MS-HuBERT introduces additional complexity to the pre-training process, particularly with the incorporation of the Swap method and Multicluster loss approach. This increased complexity may result in higher computational costs increasing the total time for pre-training only. During inference MS-HuBERT and HuBERT follows the same forward pass.
6 Conclusion
Our results highlight the potential of MS-HuBERT in bridging the performance gap between HuBERT and data2vec on the ASR Librispeech benchmark and content based tasks, ASR and PR, on the SUPERB benchmark. MS-HuBERT is aimed at mitigating the pre-training and inference mismatch in masked language modeling for learning. Building upon the HuBERT framework, MS-HuBERT incorporates two key modifications: our proposed Swap method, enabling full context access during pre-training, and the Multicluster loss approach for more effective training. Through empirical evaluation on the ASR Librispeech benchmark, MS-HuBERT demonstrates significant performance improvements over the original HuBERT model, achieving state-of-the-art results and matching the performance of data2vec in high-resource settings. Future research could explore further enhancements to the MS-HuBERT methodology to avoid iterative pre-training or improving the quality pseudo labels altogether. Lastly, scaling the model size is also an open question.
7 ACKNOWLEDGMENTS
Hemant Yadav is supported by Microsoft Research India PhD Fellowship program. Rajiv Ratn Shah is partly supported by the Infosys Center for AI, the Center of Design and New Media, and the Center of Excellence in Healthcare at IIIT Delhi.

### v4

@@ Abstract
In recent years, self-supervised pre-training methods have gained significant traction in learning and encoding high-level information from speech data. Among these methods, HuBERT has demonstrated state-of-the-art performance in automatic speech recognition (ASR). However, HuBERT’s performance lags behind data2vec due to disparities in pre-training strategies. In this paper, we propose MulticlusterSwap-HuBERT (MS-HuBERT), which integrates a Swap method to address pre-training and inference mismatch observed in HuBERT and incorporates Multicluster masked prediction loss for more effective utilization of the models capacity. MS-HuBERT, an end-to-end self-supervised pre-training method for robust speech representation learning, beat vanilla HuBERT on the ASR Librispeech benchmark by a large margin. Additionally, we demonstrate that the embeddings obtained during pre-training encode essential information for improving ASR performance. The model is available to use in SUPERB repository  [^1] https://github.com/s3prl/s3prl.
@@ 1 Introduction
In the recent years, there has been a significant interest in studying self-supervised pre-training methods to learn/encode high level information present in the speech data [1, 2, 3, 4, 5]. These SSL methods utilize the input data itself to learn to encode useful information, with the choice of pretext task playing a pivotal role in the encoded information. The most popular pretext task used is masked predictive coding (MPC) [6, 1, 7, 8, 9]. HuBERT [2] is one such model that popularized the masked language modelling (MLM) technique to learn high-level speech representations from raw audio by achieving SOTA on the automatic speech recognition (ASR) task. The underlying concept of HuBERT revolves around iterative pre-training: starting with a raw audio/pseudo-label pair (x/y), the model undergoes successive training iterations where the trained model updates the pseudo-labels, iteratively refining its representations until a predefined stopping criterion is reached. I However, despite its success, HuBERT falls short compared to data2vec [10] in ASR performance for two primary reasons: firstly, during pre-training, data2vec accesses the full context to generate continuous labels, which are updated after each gradient update step, as opposed to the fixed discrete labels utilized in HuBERT for the each iteration; and secondly, the output is averaged from multiple layers for loss calculation.
@@
To bridge this gap, we propose two modifications to the HuBERT framework. Firstly, we introduce the ”Swap” method to enable full context access during pre-training, thus addressing the pre-training and inference mismatch observed in HuBERT and other MLM-based methods by using both the masked and unmasked views during pre-training. Swap is motivated by a simple idea, used heavily in the field of computer vision [11, 12, 13]. Where two augmented views of the input are used to learn a high level representation. Given $e$ layers in a encoder, it is a general practice to add a similarity loss on the output embeddings of the encoder after each layer, as seen in works using U-net type architectures [14, 15, 16]. In contrast, we propose a Swap method which swaps the output embeddings, at certain indices, after each layer of the encoder between the masked and unmasked view of the input. This is motivated by the simple fact, that the learned model is expected to generate exactly the same output regardless of the two views.
@@
Secondly, inspired by the work of Yadav et al. [17], we adopt a Multicluster masked prediction loss (MPL) approach. Using multiple cluster centers, also called multiple resolutions, has been investigated by [18, 19, 17]. In [19], the author introduces down-sampling and up-sampling modules within the transformer encoder after each layer to facilitate learning features at multiple resolutions. On the other hand, [18] explores parallel and hierarchical variations of HuBERT with findings indicating the superiority of the hierarchical approach. This involves training multiple models, each model adds almost same parameters as the original HuBERT, at various resolutions using CNN as a down-sampling module. Lastly [17] uses the fact that MPL is applied at multiple layers of encoder at different resolutions. This method does not introduce any additional parameters to the original HuBERT model, except the linear layers used for loss calculation which are discarded after the pre-training. In this work, we adopt this approach and modify it for our use case for loss calculation.
@@
Based on these observations. In this work, we propose MulticlusterSwap-HuBERT (MS-HuBERT) method, which incorporates (i) the Swap method to address the pre-training and inference mismatch issue, as the [MASK] symbol never appears during the inference, and (ii) the Multicluster MPL similar to [17]. Our contributions are as follows:
@@
    We show that MS-HuBERT outperforms the original HuBERT on the ASR Librispeech benchmark with a big margin. And matches the performance of data2vec in high-resource setting.
@@
    We showcase that the embeddings acquired during pre-training encode crucial information essential for addressing the ASR task. Thus utilizes the model capacity very effectively.
@@ 4 Results
Table 2: SUPERB fine-tuning results. P and F stand for pre-training and fine-tuning respectively. 6-11 means using layers only 6,7,8,9,10,11. For more detailed results using different layers see Section 4.4.
Method P/F PR ASR
wav2vec 2.0 [1] P 5.74 6.43
HuBERT [2] P 5.41 6.42
HuBERT + Spin256 [23] P + F 4.39 6.34
HuBERT + LASER [24] P + F (1/3) 4.61 6.18
WavLM [9] P 4.84 6.21
WavLM + Spin256 [23] P + F 4.18 6.34
WavLM + LASER [24] P + F (1/3) 4.28 5.92
WavLM Base + (iter 3) (90k) [9] P 3.92 5.59
ContentVec [25] P + F 4.9 5.7
data2vec [10] P 4.69 4.94
MR-HuBERT [19] (iter 3) P 4.16 5.76
MS-HuBERT P 4.42 5.60
-Swap P 4.37 5.68
-Multi P 5.0 6.48
MS-HuBERT (iter 3) P 4.17 5.32
MS-HuBERT (iter 3) (6 - 11) P 4.05 5.25
4.1 Main Results: Supervised Fine-tuning and Inference
@@
Notably, upon the removal of the Swap concept, we observed a degradation in performance, particularly in low-resource settings. This proves that the Swap method does indeed contribute positively to the performance gains.
@@ 4.2 MS-HuBERT as a Feature Extractor
Based on the above comparison, we hypothesize, that MPL using pseudo labels generated from k-means is better suited for PR task than ASR. The reason might be the clustering algorithm used itself.
Figure 2: Solid lines show the CCA similarity with the word labels. Dotted lines show the AUC area under the curv for different models.
Figure 3: CCA similarity with the word labels for MS-HuBERT and its variants. The S-HuBERT curve is similar to WavLM.
CCA Similarity with Word Labels: Following the layer-wise analysis conducted by Pasad et al. [27, 28], we use a modified version of canonical correlation analysis (CCA). Specifically, a projection-weighted CCA (PWCCA) [29]. The plots are shown in Figure 2. It is clear that MS-HuBERT significantly enhances the performance of word-level information across the transformer encoder layers. Additionally, we compute the area under the curve (AUC) and observe that it consistently surpasses that of Hubert. This increases the model capacity utilization compared to HuBERT. We also plot the M-HuBERT (MS-HuBERT - Swap) to study the effect the of Swap method. We found that combining Swap with the Multicluster loss slightly increases the AUC compared to not using it as shown in Figure 3.
Table 3: 3rd iteration results using a 4-gram language model. WavLM + is model trained on 90k hours of dataset with data augmentation and 1 million steps.
Method test-clean test-other
1hr
WavLM Base + (iter 3) (90k) [9] 5.4 9.8
MR-HuBERT [19] (iter 3) 7.41 12.14
Ms-HuBERT 5.9 11.3
Ms-HuBERT (iter 3) 5.4 10.8
10hr
WavLM Base + (iter 3) (90k) 4.2 8.8
MR-HuBERT [19] (iter 3) 4.91 8.33
Ms-HuBERT 4.1 8.8
Ms-HuBERT (iter 3) 3.9 8.4
100hr
WavLM Base + (iter 3) (90k) 2.9 6.8
MR-HuBERT [19] (iter 3) 3.57 6.81
Ms-HuBERT 3.0 7.1
Ms-HuBERT (iter 3) 3.0 7.0
Table 4: No LM. on the 100hr subset. Encoder-fixed. Encoder-decoder models comparison.
Method test-clean test-other
SpeechT5 [30] 4.4 10.4
speech2c [31] 5.0 11.9
MS-HuBERT (iter 3) 4.2 10.2
4.3 3rd Iteration Models
We compare 3rd iteration MS-HuBERT (iter 3) to the 3rd iteration WavLM base+ which is trained on 960 and 94,000 hours of dataset respectively. WavLM is trained for 1 million steps. 3rd iteration MS-HuBERT is trained using the six set pseudo labels generated from the 7th layer of 2nd iteration MS-HuBERT. As shown in Table 3, MS-HuBERT (iter 3) achieves comparable performance to WavLM base +, even though pre-trained using 100 times less data. Which again shows that MS-HuBERT utilizes the model capacity most effectively.
Table 4 presents our results for an encoder-decoder framework inspired by Speech2C [31]. The encoder remains fixed, while only the decoder is trained using six hierarchical clusters, each applied sequentially across the six decoder layers, with fewer clusters assigned to the initial layer. Notably, MS-HuBERT achieves superior performance.
Table 5: Evaluation of Individual Layers on the SUPERB Benchmark. PR and ASR tasks are trained for 20% of the total steps, with results in circular brackets indicating training for 100% of the total steps. Results in rectangular brackets are for HuBERT [32].
Layer no/s PR ASR
6 - 11 (4.05) (5.25)
5 8.41 [21.05] 10.52
6 6.56 8.75
7 5.63 [11.23] 7.24
8 5.01 6.24
9 4.92 [7.80] 5.62
10 4.41 (4.36) 5.21
11 4.60 [6.10] 5.04 (4.85)
12 4.82 5.17
4.4 Evaluation of Individual Layers on the SUPERB Benchmark
Rather than computing a weighted average across all layers, we analyze the performance of using a single layer with a high CCA score, specifically from layer 5 onward. The results are presented in Table 5 and Figures 4 and 5. In summary, We found that a weighted average yields the best performance for the PR task, while using a specific layer works best for the ASR task. Overall depth is important for learning better representations.
Lastly, using the 8th layer of MS-HuBERT improves over HuBERT, resulting in 33% less compute of transformer encoder.
Figure 4: PER values are plotted with the x-axis representing layers 5 to 12, ordered from left to right.
Figure 5: WER values on the y-axis should be divided by 100 to obtain the final WER. The x-axis represents layers from 5 to 12, ordered from left to right.
@@ 5 Discussion
Given MS-HuBERT is trained using the pseudo labels generated from the the first iteration HuBERT, there could be a scope for performance improvements as is evident from the MS-HuBERT (iter 3). Or the number of layers used in between the Multicluster MPL may restrict the capacity to learn higher level embeddings. However, larger models may alleviate this limitation by offering more layers in between the MPL calculation.
6 Conclusion and Future Work
Our results highlight the potential of MS-HuBERT in bridging the performance gap between HuBERT and data2vec on the ASR Librispeech benchmark and content based tasks, ASR and PR, on the SUPERB benchmark. MS-HuBERT is aimed at mitigating the pre-training and inference mismatch in masked language modeling for learning. Building upon the HuBERT framework, MS-HuBERT incorporates two key modifications: the Swap method, enabling full context access during pre-training, and the Multicluster loss approach for more effective training. Through empirical evaluation on the ASR Librispeech benchmark, MS-HuBERT demonstrates significant performance improvements over the original HuBERT model, achieving state-of-the-art results and matching the performance of data2vec in high-resource settings. Future research could explore further enhancements to the MS-HuBERT methodology to avoid iterative pre-training or improving the quality pseudo labels altogether. Lastly, scaling the model size is also an open question.
7 Limitations
Complexity and Computational Cost: MS-HuBERT introduces additional complexity to the pre-training process, particularly with the incorporation of the Swap method and a little overhead when using Multicluster MPL. This increased complexity may result in higher computational costs during pre-training only. But during inference, MS-HuBERT and HuBERT has the same complexity as they share exactly the same architecture.
Appendix A appendix
A.1 CCA
Figure 6, 7 and, 8 shows the plots for CCA similarity for different settings similar to the original work in [28, 27].
A.2 MS-HuBERT components
Table 6 shows the ablation results for different components of MS-HUBERT (iter 2).
Figure 6: CCA similarity with the mel
Figure 7: CCA similarity with the phone
Figure 8: CCA similarity with the word
Table 6: ASR ablation
Method LM test-lean test-other
1hr
MS-HuBERT None 18.4 24.8
M-HuBERT None 19.5 26.3
S-HuBERT None 21.3 28.2
MS-HuBERT 4-gram 5.9 11.3
M-HuBERT 4-gram 6.2 12.3
S-HuBERT 4-gram 6.1 12.2
10hr
MS-HuBERT None 8.2 14.7
M-HuBERT None 8.5 14.9
S-HuBERT None 9.7 17.0
MS-HuBERT 4-gram 4.1 8.8
M-HuBERT 4-gram 4.1 9.2
S-HuBERT 4-gram 4.3 9.5
100hr
MS-HuBERT None 4.6 10.6
M-HuBERT None 4.8 10.9
S-HuBERT None 5.5 12.5
MS-HuBERT 4-gram 3.0 7.2
M-HuBERT 4-gram 3.1 7.4
S-HuBERT 4-gram 3.2 7.9